OpenStudy (anonymous):

if m,n equal to or greater than 2 are integers, find the critical points of f(x) x^m (1-x)^n there are 3 critical points what are they? HELP

OpenStudy (anonymous):

\[f(x)=x^m(1-x)^n~~\Rightarrow~~f'(x)=mx^{m-1}(1-x)^n-nx^m(1-x)^{n-1}\] When is this equal to zero? \[\begin{align*}mx^{m-1}(1-x)^n-nx^m(1-x)^{n-1}&=0\\ x^m(1-x)^n\left(\frac{m}{x}-\frac{n}{1-x}\right)&=0 \end{align*}\] Two of the solutions are readily apparent, \(x=0\) and \(x=1\). \[\begin{align*}\frac{m}{x}-\frac{n}{1-x}&=0\\ \frac{m(1-x)-nx}{x(1-x)}&=0&\text{Presumably, }x\text{ is neither 0 nor 1.}\\ m(1-x)-nx&=0\\ m-(m+n)x&=0\\ m&=(m+n)x\\ x&=\frac{m}{m+n} \end{align*}\]

OpenStudy (anonymous):

See the graph for m=2 and n=3 http://www.wolframalpha.com/input/?i=plot++%281+-+x%29%5E3+x%5E2+from+x%3D-.1+to+x%3D1.4 You cansee the 3 critical points

OpenStudy (anonymous):

thank you